OpenStudy (anonymous):

A punter kicks a football upward with an initial speed of 48 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (anonymous):

So it looks like this? \[h = 48t-16t^2\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so now what?

OpenStudy (anonymous):

Graph it

OpenStudy (anonymous):

cant graph that

OpenStudy (anonymous):

Your h is 36

OpenStudy (campbell_st):

so you need to le the height = 0 and then solve 0 = 48t - 16t^2 for t this can be factored

OpenStudy (anonymous):

Now try to put it into quadratic form like \[48t-16t^2-36=0\]

OpenStudy (anonymous):

Then factor this and you should get 2 values.

OpenStudy (anonymous):

4(8t-9)

OpenStudy (anonymous):

No...use quadratic equation

OpenStudy (anonymous):

t=3,0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats the amswer?

OpenStudy (anonymous):

so the answer is 3 seconds

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

A company launches 4 new products. The market price, in dollars, of the 4 products after a different number of years, x, is shown in the following table: Product Function Year 1 (dollars) Year 2 (dollars) Year 3 (dollars) Product 1 f(x) = x3 1 8 27 Product 2 g(x) = x2 + 11 12 15 20 Product 3 h(x) = 4x 4 16 64 Product 4 j(x) = 4x + 8 12 16 20 Based on the data in the table, for which product does the price eventually exceed all others? Product 1 Product 2 Product 3 Product 4

OpenStudy (anonymous):

can u help with this

OpenStudy (anonymous):

h(x) is this? \[4x^4\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

If it is product 3 will exceed all other price, it's increasing exponentially

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

It's product 1, product 3 is only 4x

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, f(t), in feet, at different times, t, in seconds: f(t) = −16t2 + 48t + 100 The average rate of change of f(t) from t = 3 seconds to t = 5 seconds is _____feet per second.

OpenStudy (anonymous):

\[\left[\begin{matrix}Product & Function & Year 1 & Year 2 & Year 3\\ 1 & f(x)=x^3 & 1 &8 &27 \\ 2 & g(x) = x^2+11 & 12 &15&20 \\ 3 & h(x) = 4x & 4&16& 64 \\4 &j(x)=4x+8 &12 &16 &20\end{matrix}\right]\]

OpenStudy (anonymous):

that should seem clearer when i put it this way

OpenStudy (anonymous):

Ok take the derivative of the function f(t)

OpenStudy (anonymous):

I hope you know how to do that

OpenStudy (anonymous):

can i just solve for t

OpenStudy (anonymous):

no you can't solve for t because every value of t gives you something new, its a function

OpenStudy (anonymous):

and you are finding the rate of change

OpenStudy (anonymous):

This curve looks like this

OpenStudy (anonymous):

so you are solving for the tangent line like this